# Self-propulsion of floating ice blocks caused by melting in water

### v2

@@ IV Propulsion model for inclined ice blocks melting in freshwater
Figure 5: A schematic of the time-averaged current due to the thermal convection flow with velocity $\bm{v}$ develops below the ice block in a fluid layer with thickness $\delta_{v}$ which increases with distance $x^{\prime}$ from the front tip. The current leaves the control volume indicated by the dashed line below the ice block, over a length $\delta_{x}$ with a velocity $\bm{v_{p}}$, generating a horizontal thrust in the opposite direction which is balanced by fluid drag in the stationary regime. Consequently, the ice block moves in the $x$ direction with a steady velocity $\bm{U_{b}}$.
Figure 6: A schematic section of the melting ice-water interface and the current. The velocity of the ice/water interface during melting $\bm{v_{m}}=-v_{m}\,\bm{e_{y^{\prime}}}$, defines the melting velocity $v_{m}$. The graph on the left depicts the temperature evolution as a function of the distance to the block $y^{\prime}$. The thermal boundary layer is assumed to have a constant thickness $\delta_{T}$ under steady melting conditions. For simplification, we assume that the increase of the temperature between the ice temperature $T_{i}\equiv T_{m}=0\,^{\circ}$ and the bath temperature $T_{b}$ is linear. When $\delta_{T}$ becomes sufficiently thick, the layer is destabilized by gravity and emits plumes which then sink. These plumes feed a gravity current that moves along the inclined surface with velocity $\bm{v}$. Because the water density is maximal at $T=T_{c}$, we consider the instability of the layer of thickness $\delta_{i}$, which is at a temperature between $T_{c}$ and $T_{b}$ to obtain an estimate of $\delta_{T}$.
@@ IV.1 Melting driven by thermal convection
We evaluate the melting velocity $v_{m}$ of a floating inclined ice block valid when $T_{b}>T_{c}=3.98\,^{\circ}$, the temperature where freshwater density is maximum. The velocity of the receding ice/water interface during melting is directed locally perpendicular to this interface and its magnitude defines $v_{m}$ (see Fig. 6). We consider the melting as being driven by thermal convection because the block is initially stationary, and the water below the block is unstable due to gravity as it cools. Keitzl, et al. [30] investigated experimentally and numerically the melting of ice in fresh water for a flat ice roof suspended above a water bath. These authors also derived a scaling law for the melting velocity driven by thermal convection. However, their model is arbitrarily calibrated using numerical simulations. Here, we propose a model to estimate the melting velocity $v_{m}$ when the interface is inclined, and shown to be consistent with that by Keitzl, et al.  [30].
We consider an ice block with a volume which is much smaller than the water bath volume, and assume that the bath temperature $T_{b}$ is constant, sufficiently far from the block. The ice block surface temperature $T_{i}$ is assumed to be at the ice melting temperature $T_{m}=0\,^{\circ}$. The latent heat of ice melting $\mathcal{L}=333.5$ kJ kg-1, and the heat capacity of ice is $C_{p,i}=2.11$ kJ kg-1 K-1. The energy required to heat the ice block to $T_{i}$ for an initial ice block temperature of $-15\,^{\circ}$ can be thus considered to be negligible, because $\mathcal{L}/C_{p,i}\gg 15\,^{\circ}$. Further, the ice blocks are allowed to rest in our experiments for at least $\tau_{rest}=10$ minutes after removal from the freezer. As the ice diffusivity is $\kappa_{ice}=1.11\times 10^{-6}$ m2 s-1, the temperature typically diffuses over a distance $\sqrt{\kappa_{ice}\,\tau_{rest}}\approx 26$ mm. Therefore, we neglect the heat transport inside the block for ice blocks that are a few tens of centimeters in size, and assume that the inner temperature of the ice blocks near the melting interface is $T_{m}$.
@@
where $\rho_{c}=999.96$ kg m-3 is the maximal water density at $T_{c}=3.98\,^{\circ}$, and coefficient $\beta=7\times 10^{-6}$ K-2. Thus, $\rho$ of water near the ice block is greater than that of the warmer bath water when $T_{b}>T_{c}$, and this density difference drives a convection flow. The temperature dependence of $\rho$ is more accurately described by Bigg’s empirical relation [31, 32]. By fitting $\rho(T)$ given by Bigg’s relation, the coefficient $\beta=6.5\times 10^{-6}$ K-2 in Eq. (1) is obtained, and used henceforth. Under steady state conditions, we assume turbulent thermal convection, i.e. on average the temperature change is localized to a thin thermal boundary layer of thickness $\delta_{T}$, where the temperature increases from $T_{i}$ to $T_{b}$ as illustrated in Fig. 6. To simplify the modeling, we assume a linear temperature profile. The heat transport is diffusive in the boundary layer and convective outside. In the turbulent region below, the sinking plumes and upwelling flow transfer heat effectively. Due to the inclination of the block, the convection flow self-organizes into a current with characteristic velocity $\bm{v}$ on the scale of the ice block, which escapes the control volume with a velocity $\bm{v_{p}}$ which is directed below and behind the block (see Fig. 5). This directed flow is the origin of the propulsion mechanism.
@@
We consider that the melting velocity $v_{m}$ of the solid-liquid interface as being controlled by the convection flow. We denote $y^{\prime}$ as the coordinate normal to the melting interface and $y^{\prime}_{c}$ as the distance to where $T=T_{c}$. Because $\rho$ is below $\rho_{c}$ where $T_{i}<T<T_{c}$ over $0<y^{\prime}<y^{\prime}_{c}$, the fluid layer is stable relatively to gravity. In contrast, the domain $y^{\prime}_{c}<y^{\prime}<\delta_{T}$ is denser than the bath at the density $\rho(T_{b})$ and can be susceptible to convection instability. Assuming a linear temperature profile, we have $y^{\prime}_{c}=\delta_{T}\,(T_{c}-T_{i})/(T_{b}-T_{i})$ and consequently the thickness of the unstable layer is $\delta_{i}=\delta_{T}-y^{\prime}_{c}=\delta_{T}\,(T_{b}-T_{c})/(T_{b}-T_{i})$. We evaluate the dimensionless density contrast by dividing $\Delta\rho$ the density difference between the fluid at $T_{c}$ and the fluid at $T_{b}$, by the average density between these two values:
$\dfrac{\Delta\rho}{\rho}=2\Big{(}\dfrac{\rho_{c}-\rho_{b}}{\rho_{c}+\rho_{b}}\Big{)}\,,$ (2)
@@
where $\rho_{b}=\rho(T_{b})$ is computed using Eq. 1. If $\Delta\rho$ is sufficiently high, a Rayleigh-Bénard instability will occur. According to previous studies in geometries with semi-infinite extent under steady state conditions [33, 34, 35], and as discussed in a lecture by Linden [36], the thickness of the boundary layer remains on average close to the critical value at the onset of the Rayleigh-Bénard instability to first approximation. From the definition of the Rayleigh number $Ra$ and assuming the critical value of $Ra=Ra_{c}$, we have
@@
where $g\,\cos\theta$ is the projection of the gravitation acceleration perpendicular to the ice block surface and $\kappa$ is the thermal diffusivity of water. To obtain an estimate, we use the parameter values from Keitzl, et al. [30], where their temperature dependence is neglected. We take the value of $\kappa=1.33\times 10^{-7}$ m2 s-1 at $T=0\,^{\circ}$, because between $T=0\,^{\circ}$ and $T=30\,^{\circ}$, $\kappa$ changes less than $10\%$ [18]. However, the decrease in $\nu$ with temperature is more significant because $\nu=1.79\times 10^{-6}$ m2 s-1 at $T=0\,^{\circ}$, $\nu=1.58\times 10^{-6}$ m2 s-1 at $T=4\,^{\circ}$ and $\nu=1.00\times 10^{-6}$ m2 s-1 at $T=20\,^{\circ}$[18]. Accordingly, the thermal Prandtl number in water $Pr=\nu/\kappa$ is about $13.5$ at $T=0\,^{\circ}$, $11.9$ at $T=4\,^{\circ}$, and $7.5$ at $20\,^{\circ}$. To simplify, we choose the average value of $\nu$ between $T_{c}$ and $T_{b}$.
@@
after rearranging Eq. (3). For $T_{b}=20\,^{\circ}$ and $\theta=26.5^{\circ}$, $\nu=(\nu(T_{c})+\nu(T_{b}))/2=1.29\times 10^{-6}$ m2 s-1 and $\Delta\rho/\rho\approx\beta\,(T_{b}-T_{c})^{2}\approx 1.80\times 10^{-3}$, we find $\delta_{i}\approx 1.93$ mm and $\delta_{T}\approx 2.40$ mm. The experimental and numerical study by Du, et al. [38] for freezing of salt water provide correlation laws for $\nu$ and $\kappa$. With these more accurate values of the fluid properties, we find $\delta_{T}\approx 2.44$ mm, which is close to $\delta_{T}\approx 2.40$ mm obtained using Eq. (4).
[…]
@@
To find the fitting constant $\Gamma$, a set of measurements of $v_{m}$ were performed with inclined blocks. Clear ice blocks with dimensions $100\times 40\times 100$ mm3 were fixed at a prescribed inclination over the range of $\theta$ where the ice blocks were found to be stable. The displacement of the bottom interface while subjected to a detached thermal convection flow was monitored using shadowgraph imaging and backlight imaging. The measured $v_{m}$ with these two methods are shown in Fig. 7(a). While the measurement scatter is considerable, the two methods yield $v_{m}$ which are consistent with each other. Then, we estimate $\Gamma=2.2$ by fitting Eq. (7) to the data in Fig. 7(a). (Because the angle dependence is $(\cos\theta)^{1/3}$, the dependence on angle is small, and within measurement scatter.)
Figure 7: (a) The experimental measurements of $v_{m}$ at the bottom surface of an ice block as a function of its inclination in fresh water. The blocks are held at rest, $T_{b}=22\,\,^{\circ}$ and their dimensions correspond to Data Set F (clear ice). Two illumination methods are used, shadowgraph and backlight. These measurements enable us to calibrate the ice melting model in fresh water and set $\Gamma=2.2$ in Eq. (7). The theoretical estimate of Keitzl, et al. [30] after including an inclination dependency according to Eq. (8) predicts relatively well the measurements. (b) The calculated melting velocity of the ice block in freshwater with $\theta=24^{\circ}$ as a function of $T_{b}$ using Eq. (7) for $\Gamma=1$ and $\Gamma=2.2$. The theoretical estimate of Keitzl, et al. [30] $v_{m,K}$ according to Eq. (8) is also plotted. Both models are valid for $T>T_{c}$.
The melting velocity estimated by Keitzl, et al[30], after multiplying gravitational acceleration $g$ by a factor $\cos\theta$ to include the effect of the block inclination, can be written as,
@@
and $\beta\,T_{c}$ is a density contrast according to Eq. (1) between the maximal water density and the water density at $T_{m}=0\,^{\circ}$. They considered $Pr=10$ and $Ra_{c}=1000$, comparable to the values in our study, and calibrated the temperature dependency using numerical simulations of the temperature profile at the melting interface [30] while testing the predicted values only over temperatures ranging from $4.5\,^{\circ}$ to $14.8\,^{\circ}$ with experiments. Plotting $v_{m,K}$ in Fig. 7(a), we observe that it is somewhat lower compared with our experimental data, but reasonable considering the apparent experimental fluctuations. We also plot $v_{m,K}$ using Eq. (8), and our calculated $v_{m}$ using Eq. (7) with fit $\Gamma=2.2$, as well as with fit $\Gamma=1$ in Fig. 7(b) as a function of $T_{b}$ for $T_{b}>T_{c}$, where both models are valid. The two models provide similar predictions of melting velocity and the same scaling at large $T_{b}$.
Further, our calculation of $v_{m}$ (Eq. 7) gives the same scaling as previously found by Kerr in Appendix B of Ref. [40], but without taking into account the nonlinear evolution of water density with temperature. The effect of this nonlinear evolution has been investigated in numerical studies of Rayleigh-Bénard instability in presence of a upper stable stratified layer above a unstable region [41, 42, 43], but without phase change. Under standard conditions for ice melting, the dynamics of the upper stable layer may be decoupled from the thermal convection flow below [43]. As in Kerr [40] and Keitzl [30], we neglect the contribution of the melt fluid in the calculation of the melting velocity. The layer of melt fluid at $T_{m}$ thickens the stable stratified layer, insulating the ice block, and reducing the melting velocity. Using numerical simulations, Keitzl et al. [30] estimate a decrease of less than $10$% of the meting rate for bath temperature smaller than $20\,^{\circ}$. Finally, after non-dimensionalizing Eq. (7), and using a characteristic length scale $L^{\star}$ large compared to $\delta_{T}$, it can be shown that the dimensionless thermal flux, the Nusselt number $Nu$, is proportional to the Rayleigh number to the power $1/3$. In the general context of thermal convection, this scaling law corresponds to the regime where the heat flux is controlled by the thermal boundary layer [33, 44, 45]. The scaling $Nu\propto Ra^{1/3}$ has been also reported experimentally for the melting of ice cylinders about ten centimeter of diameters in fresh and saline water [26, 46].

### v3

@@ IV Propulsion model for inclined ice blocks melting in freshwater
Figure 5: Schematic of the melting driven propulsion mechanism. A time-averaged current due to the thermal convection flow with velocity $\bm{v}$ develops below the ice block in a fluid layer with thickness $\delta_{v}$ which increases with distance from the front tip. $x^{\prime}$ is the coordinate along the inclined wall from the rear to the tip of the ice block. The current leaves the control volume indicated by the dashed line below the ice block, over a length $\delta_{x}$ with a velocity $\bm{v_{p}}$, generating a horizontal thrust in the opposite direction which is balanced by fluid drag in the stationary regime. Consequently, the ice block moves in the $x$ direction with a steady velocity $\bm{U_{b}}$.
Figure 6: A schematic section of the melting ice-water interface and the current. The velocity of the ice/water interface during melting $\bm{v_{m}}=-v_{m}\,\bm{e_{y^{\prime}}}$, defines the melting velocity $v_{m}$. The graph on the left depicts the temperature evolution as a function of the distance to the block $y^{\prime}$. The thermal boundary layer is assumed to have a constant thickness $\delta_{T}$ under steady melting conditions. For simplification, we assume that the increase of the temperature between the ice temperature $T_{i}\equiv T_{m}=0\,^{\circ}$C and the bath temperature $T_{b}$ is linear. Because the water density is maximal at the temperature of maximum density $T_{c}$, the domain for which $T_{m}<T<T_{c}$ is buoyancy stable, whereas the domain for which $T_{c}<T<T_{b}$ is instable. We consider the instability of this layer of thickness $\delta_{i}$ , which emits sinking plumes. These plumes feed a gravity current that moves along the inclined surface with velocity $\bm{v}$. By linearity of the temperature profile, the estimation of $\delta_{T}$ is deduced from the one of $\delta_{i}$.
@@ IV.1 Melting driven by thermal convection
We evaluate the melting velocity $v_{m}$ of a floating inclined ice block valid when $T_{b}>T_{c}=3.98\,^{\circ}$C, the temperature where freshwater density is maximum. The velocity of the receding ice/water interface during melting is directed locally perpendicular to this interface and its magnitude defines $v_{m}$ (see Fig. 6). We consider the melting as being driven by thermal convection because the block is initially stationary, and the water below the block is unstable due to gravity as it cools. Keitzl, et al. [32] investigated experimentally and numerically the melting of ice in fresh water for a horizontal ice lid immersed in a water bath. These authors also derived a scaling law for the melting velocity driven by thermal convection. However, their model is arbitrarily calibrated using numerical simulations. Here, we propose a model to estimate the melting velocity $v_{m}$ when the interface is inclined, and shown to be consistent with that by Keitzl, et al.  [32].
We consider an ice block with a volume which is much smaller than the water bath volume, and assume that the bath temperature $T_{b}$ is constant, sufficiently far from the block. The ice block surface temperature $T_{i}$ is assumed to be at the ice melting temperature $T_{m}=0\,^{\circ}$C. The latent heat of ice melting is $\mathcal{L}=333.5$ kJ kg-1, and the heat capacity of ice is $C_{p,i}=2.11$ kJ kg-1 K-1. The energy required to heat the ice block to $T_{i}$ for an initial ice block temperature of $15\,^{\circ}$C can be thus considered to be negligible, because $\mathcal{L}/C_{p,i}\gg 15$K. Further, the ice blocks are allowed to rest in our experiments for at least $\tau_{rest}=10$ minutes after removal from the freezer. As the ice diffusivity is $\kappa_{ice}=1.11\times 10^{-6}$ m2 s-1, the temperature typically diffuses over a distance $\sqrt{\kappa_{ice}\,\tau_{rest}}\approx 26$ mm. Therefore, we neglect the heat transport inside the block for ice blocks that are a few tens of centimeters in size, and assume that the inner temperature of the ice blocks near the melting interface is $T_{m}$.
@@
where $\rho_{c}=999.96$ kg m-3 is the maximal water density at $T_{c}=3.98\,^{\circ}$C, and coefficient $\beta=7\times 10^{-6}$ K-2. Thus, $\rho$ of water near the ice block is greater than that of the warmer bath water when $T_{b}>T_{c}$, and this density difference drives a convection flow. The temperature dependence of $\rho$ is more accurately described by Bigg’s empirical relation [33, 34]. By fitting $\rho(T)$ given by Bigg’s relation, the coefficient $\beta=6.5\times 10^{-6}$ K-2 in Eq. (1) is obtained, and used henceforth. Under steady state conditions, we assume turbulent thermal convection, i.e. on average the temperature change is localized to a thin thermal boundary layer of thickness $\delta_{T}$, where the temperature increases from $T_{i}$ to $T_{b}$ as illustrated in Fig. 6. To simplify the modeling, we assume a linear temperature profile. The heat transport is diffusive in the boundary layer and convective outside. In the turbulent region below, the sinking plumes and upwelling flow transfer heat more efficiently than thermal diffusion. Due to the inclination of the block, the convection flow self-organizes into a current with characteristic velocity $\bm{v}$ on the scale of the ice block, which escapes the control volume with a velocity $\bm{v_{p}}$ which is directed below and behind the block (see Fig. 5). This directed flow is the origin of the propulsion mechanism.
@@
We consider the case where the melting velocity $v_{m}$ of the solid-liquid interface is controlled by the convection flow. We denote $y^{\prime}$ as the coordinate normal to the melting interface and $y^{\prime}_{c}$ as the distance to where $T=T_{c}$. Because $\rho$ is below $\rho_{c}$ where $T_{i}<T<T_{c}$ over $0<y^{\prime}<y^{\prime}_{c}$, the fluid layer is stable relative to a density inversion instabilities due to gravity, leading to detached flows as in the Rayleigh-Bénard or Rayleigh-Taylor instabilities. In contrast, the domain $y^{\prime}_{c}<y^{\prime}<\delta_{T}$ is denser than the bath at the density $\rho(T_{b})$ and can be susceptible to such convection instabilities. Assuming a linear temperature profile, we have $y^{\prime}_{c}=\delta_{T}\,(T_{c}-T_{i})/(T_{b}-T_{i})$ and consequently the thickness of the unstable layer is $\delta_{i}=\delta_{T}-y^{\prime}_{c}=\delta_{T}\,(T_{b}-T_{c})/(T_{b}-T_{i})$. We evaluate the dimensionless density contrast by dividing $\Delta\rho$ the density difference between the fluid at $T_{c}$ and the fluid at $T_{b}$, by the average density between these two values:
$\dfrac{\Delta\rho}{\rho}=2\left(\dfrac{\rho_{c}-\rho_{b}}{\rho_{c}+\rho_{b}}\right)\,$, (2)
@@
where $\rho_{b}=\rho(T_{b})$ is computed using Eq. 1. Large enough values of $\Delta\rho$ may trigger Rayleigh-Bénard-like instabilities. According to previous studies in geometries with semi-infinite extent under steady state conditions [35, 36, 37], and as discussed in a lecture by Linden [38], the thickness of the boundary layer remains on average close to the critical value at the onset of the Rayleigh-Bénard instability to first approximation. From the definition of the Rayleigh number $Ra$ and assuming the critical value of $Ra=Ra_{c}$, we have
@@
where $g\,\cos\theta$ is the projection of the gravitation acceleration perpendicular to the ice block surface and $\kappa$ is the thermal diffusivity of water. To obtain an estimate, we use the parameter values from Keitzl, et al. [32], where their temperature dependence is neglected. We take the value of $\kappa=1.33\times 10^{-7}$ m2 s-1 at $T=0\,^{\circ}$C, because between $T=0\,^{\circ}$C and $T=30\,^{\circ}$C, $\kappa$ changes less than $10\%$ [18]. However, the decrease in $\nu$ with temperature is more significant because $\nu=1.79\times 10^{-6}$ m2 s-1 at $T=0\,^{\circ}$C, $\nu=1.58\times 10^{-6}$ m2 s-1 at $T=4\,^{\circ}$C and $\nu=1.00\times 10^{-6}$ m2 s-1 at $T=20\,^{\circ}$[18]. Accordingly, the thermal Prandtl number in water $Pr=\nu/\kappa$ is about $13.5$ at $T=0\,^{\circ}$C, $11.9$ at $T=4\,^{\circ}$C, and $7.5$ at $20\,^{\circ}$C. To simplify, we choose the average value of $\nu$ between $T_{c}$ and $T_{b}$.
@@
after rearranging Eq. (3). For $T_{b}=20\,^{\circ}$C and $\theta=26.5^{\circ}$, $\nu=(\nu(T_{c})+\nu(T_{b}))/2=1.29\times 10^{-6}$ m2 s-1 and $\Delta\rho/\rho\approx\beta\,(T_{b}-T_{c})^{2}\approx 1.80\times 10^{-3}$, we find $\delta_{i}\approx 1.93$ mm and $\delta_{T}\approx 2.40$ mm. The experimental and numerical study by Du, et al. [40] for freezing of salt water provide correlation laws for $\nu$ and $\kappa$. With these more accurate values of the fluid properties, we find $\delta_{T}\approx 2.44$ mm, which is close to $\delta_{T}\approx 2.40$ mm obtained using Eq. (4).
The wavelength at the marginal instability is approximately $\delta_{i} [3$9], which gives a typical plume size of $5.8$ mm. This spacing between plumes is not visible in shadowgraph images of the inclined blocks because the flow shows a complex pattern of entangled plumes. However, it can be seen more clearly for horizontal blocks (see Movie S3 [21]). In this case, we find an average distance between plumes in steady convection regime of about $5$ mm, which is close to the wavelength expected for the marginal instability. This characteristic length for melting plumes is significantly larger than for the plumes caused by dissolution of salt or sugar in water, which are approximately 0.33 mm and 1 mm, respectively [37, 41, 9].
@@
To find the fitting constant $\Gamma$, a set of measurements of $v_{m}$ were performed with inclined blocks. Clear ice blocks with dimensions $100\times 40\times 100$ mm3 were fixed at a prescribed inclination over the range of $\theta$ where the ice blocks were found to be stable. The displacement of the bottom interface while subjected to a detached thermal convection flow was monitored using shadowgraph imaging and backlight imaging. The measured melting velocities $v_{m}$ with these two methods are shown in Fig. 7(a). While the measurement scatter is considerable, the two methods yield $v_{m}$ which are consistent with each other. Then, we estimate $\Gamma=2.2$ by fitting Eq. (7) to the data in Fig. 7(a). Because the angle dependence is $(\cos\theta)^{1/3}$, the dependence on angle is small, and within measurement scatter.
[…]
The melting velocity estimated by Keitzl, et al[32] $v_{m,K}$, after multiplying gravitational acceleration $g$ by a factor $\cos\theta$ to include the effect of the block inclination, can be written as,
@@
and $\beta\,T_{c}$ is a density contrast according to Eq. (1) between the maximal water density and the water density at $T_{m}=0\,^{\circ}$C. They considered $Pr=10$ and $Ra_{c}=1000$, comparable to the values in our study, and calibrated the temperature dependency using numerical simulations of the temperature profile at the melting interface [32] while testing the predicted values only over temperatures ranging from $4.5\,^{\circ}$C to $14.8\,^{\circ}$C with experiments. Plotting $v_{m,K}$ in Fig. 7(a), we observe that it is somewhat lower compared with our experimental data, but reasonable considering the apparent experimental fluctuations. We also plot $v_{m,K}$ using Eq. (8), and our calculated $v_{m}$ using Eq. (7) with fit $\Gamma=2.2$, as well as with fit $\Gamma=1$ in Fig. 7(b) as a function of the bath temperature $T_{b}$ for $T_{b}>T_{c}$, where both models are valid. The two models provide similar predictions of melting velocity and the same scaling at large $T_{b}$.
Further, our calculation of $v_{m}$ (Eq. 7) gives the same scaling as previously found by Kerr in Appendix B of Ref. [42], but without taking into account the nonlinear evolution of water density with temperature. The effect of this nonlinear evolution has been investigated in numerical studies of Rayleigh-Bénard instability in presence of a upper stable stratified layer above a unstable region [43, 44, 45], but without phase change. Under standard conditions for ice melting, the dynamics of the upper stable layer may be decoupled from the thermal convection flow below [45]. As in Kerr [42] and Keitzl [32], we neglect the contribution of the melt fluid in the calculation of the melting velocity. The layer of melt fluid at $T_{m}$ thickens the stable stratified layer, insulating the ice block, and reducing the melting velocity. Using numerical simulations, Keitzl et al. [32] estimate a decrease of less than 1010% of the meting rate for bath temperature smaller than $20\,^{\circ}$C. Finally, after non-dimensionalizing Eq. (7), and using a characteristic length scale $L^{\star}$ large compared to $\delta_{T}$, it can be shown that the dimensionless thermal flux, the Nusselt number $Nu$, is proportional to the Rayleigh number to the power $1/3$. In the general context of thermal convection, this scaling law corresponds to the regime where the heat flux is controlled by the thermal boundary layer [35, 46, 47]. The scaling $Nu\propto Ra^{1/3}$ has been also reported experimentally for the melting of ice cylinders about ten centimeter of diameters in fresh and saline water [26, 48].